the point of intersection of the pair of the pair of e lines represented by xy+x+y+1=0 is

  1.  the point of intersection of the pair of the pair of elines represented by xy+x+y+1=0 is

Grade:9

1 Answers

Harsh Patodia IIT Roorkee
8 years ago
xy + x + y + 1 =0
x(y+1) + 1(y+1)=0
(x+1)(y+1)=0

point is (-1,-1)